Find the Tebnan equivalent circuit from terminal a,b in the next circuit.
4A 100 UR 00V U OP b
Answer #1

Find the Thevenin Equivalent Circuit dia a ton 4015 Cov 21 CB 7 + Vth ob 2=OA (2=UA apply KVL ; * -60410(1,-12)+406,-3) -0 50finding Pitha opened 8 10 ME 4A boy shorted Roth 40 * Rth= (olluo)+8 (10x40 voryo Rth = 8+8 = 168 (Rthlon Thevenin Equivalent